OpenStudy (milliedelongg):

Suppose that ​$30,000 is invested at 9​% interest. Find the amount of money in the account after 5 years if the interest is compounded annually.

OpenStudy (milliedelongg):

need help setting up the equation :c

jimthompson5910 (jim_thompson5910):

The compound interest formula is \[\LARGE A = P*\left(1+\frac{r}{n}\right)^{n*t}\] where, A = final amount in account after t years P = amount deposited r = interest rate n = compound frequency (how many times you compound interest per year) t = time in years

OpenStudy (milliedelongg):

so how would n work? you would compound it 12 times correct? or wait is it 5?

OpenStudy (danjs):

After every year, the value of the account will increase by 9%, multiply by 1.09, so you can take the initial amount at t=0 years, P=30000, and it forms the exponential A = P*(1.09)^t

OpenStudy (milliedelongg):

ohh I think 5, right?

jimthompson5910 (jim_thompson5910):

`the interest is compounded annually` so you compound 1 time per year @milliedelongg so n = 1

OpenStudy (milliedelongg):

but it says five years. does that matter? would it still be one?

jimthompson5910 (jim_thompson5910):

after plugging in n = 1 and r = 0.09, it will simplify to what @DanJS wrote out After that, plug in t = 5 because we want the amount after 5 years

jimthompson5910 (jim_thompson5910):

or you can just plug in everything at once. It's up to you

OpenStudy (milliedelongg):

ohh okay I understand

OpenStudy (milliedelongg):

I got 46,158.72 (rounded to nearest cent)

jimthompson5910 (jim_thompson5910):

same here

jimthompson5910 (jim_thompson5910):

nice work

OpenStudy (milliedelongg):

okay coolio thank you so much!!!!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (milliedelongg):

can you help me with another one? this one doesn't give me a N

jimthompson5910 (jim_thompson5910):

sure go ahead

OpenStudy (milliedelongg):

5000 is invested at 4% "continuously" for 7 years

jimthompson5910 (jim_thompson5910):

the keyword "continuously" means n essentially heads off to infinity. Basically you compound every chance you get, ie, as much as possible You'll use this formula now \[\LARGE A = P*e^{r*t}\] the 'e' is a constant that is a lot like pi = 3.1415... e = 2.71828182846 and that decimal expansion goes on forever without a known pattern

jimthompson5910 (jim_thompson5910):

to get the best accuracy, it's best to use the built in "e" button on your calculator

OpenStudy (milliedelongg):

I got 5020.49

jimthompson5910 (jim_thompson5910):

what calculator do you have?

OpenStudy (milliedelongg):

ti-84

jimthompson5910 (jim_thompson5910):

did you type the 2ND button and then the LN button to get the 'e' ?

OpenStudy (milliedelongg):

ohhh that e has a exponent thing

jimthompson5910 (jim_thompson5910):

yes tell me what you get when you retry

OpenStudy (milliedelongg):

I still got the same answer. what are you getting?

jimthompson5910 (jim_thompson5910):

you should have 5000*e^(0.04*7) typed in

OpenStudy (milliedelongg):

ohh I had it in wrong

OpenStudy (danjs):

do you remember how to show the limit used here for e in another tab when you free?